# Intersection and Union Hierarchies of Deterministic

### v2

@@ 1.2 New Separation Results
As noted in Section 1.1, some of the fundamental properties associated with $\mathrm{DCFL}(d)$ heavily rely on the single separation result (*) of Liu and Weiner. However, Liu and Weiner’s technical tool that leads to their main result does not seem to withstand a wide variety of direct applications. It is thus desirable to develop a new, simple, and practical technical tool that can find numerous applications for conducting a future study on $\mathrm{DCFL}(d)$ as well as $\mathrm{DCFL}[d]$. Our main purpose of this exposition is therefore to present a simple but powerful and practical technical tool, which is applicable to numerous languages. In a course of analyzing parse trees of (strict) dcf languages, Harrison and Havel [8] developed an iteration theorem for $\mathrm{DCFL}$. A similar approach has been taken to introduce numerous forms of pumping lemmas and iteration theorems for variants of dcf languages in, e.g., [7, 11, 12, 13, 21, 34]. Along this long line of study, this exposition also aims at developing pumping lemmas for $\mathrm{DCFL}[d]$ for any $d\geq 1$ (formulated as Lemmas 2.1 and 2.2 in Section 2.2), which may contribute to the deepening of our understandings of both $\mathrm{DCFL}[d]$ and $\mathrm{DCFL}(d)$.
@@ 2.3 Pushdown Automata in an Ideal Shape
It was shown in [27, 29] that any 1ppda can be converted into its “equivalent” 1ppda in an ideal shape. We say that two 1ppda’s are error-equivalent if, for any input $x$, their acceptance/rejection coincide.
@@ 3 New Technical Tools: Pumping Lemmas for DCFL[$d$]
In the past literature, numerous pumping lemmas (as well as iteration theorems) have been proposed. Interestingly, the form of these pumping lemmas varies. For instance, the iteration theorem of Harrison and Havel [8, Theorem 2.3] for $\mathrm{DCFL}$ focuses on a single input string in a target language, whereas Yu’s pumping lemma for $\mathrm{DCFL}$ [34, Lemma 1] deals with two input strings in the language. We wish to take the latter approach in this exposition.
@@
As a special case of $d=1$ in Lemma 3.1, we instantly obtain Yu’s pumping lemma for $\mathrm{DCFL}$ [34, Lemma 1], which was proven by a grammar-based argument. By sharp contrast, a few machine-based arguments were employed in the past literature in order to prove generic properties of $\mathrm{CFL}$ and $\mathrm{DCFL}$. As a recent example, the swapping lemma for context-free languages was proven in [24, Lemma 4.1] and [26, Corollary 4.2] by exploring various properties of one-way nondeterministic pushdown automata (or 1npda’s). Such machine-based arguments as well as various properties of machines that these arguments have brought in are not only interesting in its own right but also quite useful to analyze families of languages recognized by different types of machines, such as one-way depth-bounded storage automata [30, 31, 33], which naturally expand 1npda’s and 1dpda’s but no grammar characterizations of them are currently known to exploit.
@@ 4 Proofs of Three Separation Claims
(1) It is well-known that $Pal$ is in $\mathrm{CFL}$. Hence, we wish to verify that $Pal$ is also in $\mathrm{co}\mbox{-}\mathrm{CFL}$. Let $ODD=\{w\in\{0,1\}^{*}\mid|w|\text{ is odd }\}$. Note that $\overline{Pal}$ coincides with $\{xy\mid|x|=|y|,x,y\in\{0,1\}^{*},y\neq x^{R}\}\cup ODD$. Consider the following 1npda. Given an input $w$, nondeterministically choose either 0 or $1$. If 0 is chosen, then we check whether $w\in ODD$. Otherwise, we split $w$ nondeterministically into $xy$, store $x$ into a stack, pop $x$ symbol by symbol while reading $y$, and check that $y=x^{R}$. If $|x|\neq|y|$, then we reject $w$. Assume otherwise. If we discover any discrepancy between $y$ and $x^{R}$, then we accept $w$; otherwise, we reject it. It is obvious that this machine correctly recognizes $\overline{Pal}$, and thus $Pal$ belongs to $\mathrm{co}\mbox{-}\mathrm{CFL}$.
@@
In the case where the condition (1) fails, let us consider three factorizations $x^{\prime}=x_{1}x_{2}x_{3}$, $y=y_{1}y_{2}y_{3}$, and $z=z_{1}z_{2}z_{3}$ with $|x_{2}|\geq 1$ and $|x_{2}x_{3}|\leq c$ satisfying $x_{1}x_{2}^{i}x_{3}y_{1}y_{2}^{i}y_{3},x_{1}x_{2}^{i}x_{3}z_{1}z_{2}^{i}z_{3}\in Pal$ for any number $i\geq 0$. We conclude that $x_{2}$ is in the $(2j_{2}+2d+1)$th block because of $|x_{2}x_{3}|\leq c$. By taking $i=0$, the size of this block becomes less than $n$. Since this block is located in the right side of the center of $x^{\prime}y$, we obtain $x_{1}x_{3}y_{1}y_{3}\notin Pal$, a contradiction.
@@
The second pumping lemma then ensures the existence of two indices $j_{1},j_{2}\in[d+1]$ with $j_{1}<j_{2}$ such that $(h_{j_{1}},h_{j_{2}})$ satisfies one of the conditions (1)–(5) of the lemma. We then define $x^{\prime}=w_{1}\#\cdots w_{d+1}\#^{2}v_{1}\#\cdots\#(v_{j_{1}})^{(-)}_{\text{first}}$, $y=(\hat{v}_{j_{1}})^{(-)}_{\text{second}}\#v_{j_{1}+1}\#\cdots v_{d+1}$, and $z=(v_{j_{1}})^{(-)}_{\text{second}}\#v_{j_{1}+1}\#\cdots\#v_{j_{2}-1}\#\hat{v}_{j_{2}}\#v_{j_{2}+1}\#\cdots v_{d+1}$. Note that $h_{j_{1}}=x^{\prime}y$ and $h_{j_{2}}=x^{\prime}z$. By the above statement (*), the conditions (1)–(4) do not hold. We thus look into the condition (5). A nondegenerate iterative pair $(x_{4}x_{5},x_{7}u_{1})$ of $x^{\prime}y$ for $L^{\prime\prime}$ in the condition (5) must satisfy that $x_{4}x_{5}\sqsubseteq w_{j_{1}}$ and $x_{7}u_{1}\sqsubseteq\hat{v}_{j_{1}}$. Similarly, if a nondegenerate iterative pair $(x_{3}x_{4},u^{\prime}_{2})$ of $x^{\prime}z$ for $L^{\prime\prime}$ exists, then we obtain $x_{3}x_{4}\sqsubseteq w_{j_{2}}$ and $u^{\prime}_{2}\sqsubseteq\hat{v}_{j_{2}}$. However, this contradicts the assumption of $j_{1}<j_{2}$. The case of $(x_{2},u^{\prime}_{2})$ in the condition (5) is similarly treated. Therefore, $L^{\prime\prime}$ is not in $\mathrm{DCFL}[d]$, as requested. This completes the proof. ∎
@@
Proposition 1.5. (rephrased) For any $k\geq 2$, $k\mbox{-}\mathrm{LDA}\cap\mathrm{DCFL}[2^{k-1}]\nsubseteq(k-1)\mbox{-}\mathrm{LDA}\cup\mathrm{DCFL}[2^{k-1}-1]$.
@@
In this proof, we intend to use a series of special languages used in [27, arXiv version], each of which is in $k\mbox{-}\mathrm{LDA}$ but not in $\mathrm{DCFL}[k-1]$ for each $k\geq 2$. Let us first recall from [27, arXiv version] the language $L_{k}$, which is a slight modification of Hibbard’s [9, Section 4] original example language.
Let $k\geq 2$ be any integer. For each fixed index $i\in[k]$, we write $w_{i}$ for an arbitrary string of the form $a^{n_{i}}b^{m_{i}}c^{p_{i}}$ for any numbers $n_{i},m_{i},p_{i}\in\mathbb{N}$. The language $L_{k}$ is composed of all strings $w$ of the form $w_{2}\#w_{4}\#\cdots\#w_{k-2}\#w_{k}\#w_{k-1}\#\cdots\#w_{5}\#w_{3}\#w_{1}$ if $k$ is even, and $w_{2}\#w_{4}\#\cdots\#w_{k-1}\#w_{k}\#w_{k-2}\#\cdots\#w_{5}\#w_{3}\#w_{1}$ if $k$ is odd, together with the following conditions (a)–(c).
@@
    Let $j$ denote any index in $2,k]_{\mathbb{Z}}$. If $n_{j}\leq m_{j}$, then either $n_{j-1}=m_{j-1}$ or $m_{j-1}=p_{j-1}$ holds. If $m_{j}\leq p_{j}$, then either $n_{j-1}<m_{j-1}$ or $m_{j-1}<p_{j-1}$ holds.
@@
For instance, $L_{2}$ equals $\{a^{n_{2}}b^{m_{2}}c^{p_{2}}\#a^{n_{1}}b^{m_{1}}c^{p_{1}}\mid(n_{1}=m_{1}\wedge n_{2}=m_{2})\text{ or }(n_{1}<m_{1}\wedge m_{2}=p_{2})\}$. For convenience, we also define another language $L^{\prime}_{k}$ obtained from $L_{k}$ by replacing the last term $w_{1}=a^{n_{1}}b^{m_{1}}c^{p_{1}}$ of $L_{k}$ with $a^{p_{1}}b^{m_{1}}c^{n_{1}}$.
@@
As noted in [27, arXiv version], $L_{k}$ belongs to $k\mbox{-}\mathrm{LDA}$ but it is not in $(k-1)\mbox{-}\mathrm{LDA}$. It thus suffices to show by induction on $k$ that $L_{k}\in\mathrm{DCFL}[2^{k-1}]$ and $L_{k}\notin\mathrm{DCFL}[2^{k-1}-1]$. To do so, we want to take the following two steps (1)–(2). For readability, we assume in (1)–(2) that $n_{i}$, $m_{i}$, and $p_{i}$ express numbers in $\mathbb{N}$ for all $i$’s.
@@
(1) As the first step, we prove the membership of $L_{k}$ to $\mathrm{DCFL}[2^{k-1}]$.
@@
(i) When $k=2$, we define $A_{(1)}=\{a^{n_{2}}b^{m_{2}}c^{p_{2}}\#a^{n_{1}}b^{m_{1}}c^{p_{1}}\mid n_{1}=m_{1},n_{2}=m_{2}\}$ and $A_{(2)}=\{a^{n_{2}}b^{m_{2}}c^{p_{2}}\#a^{n_{1}}b^{m_{1}}c^{p_{1}}\mid n_{1}<m_{1},m_{2}=p_{2}\}$. Clearly, $L_{2}$ equals $A_{(1)}\cup A_{(2)}$, and thus $L_{2}$ belongs to $\mathrm{DCFL}[2]$. Similarly, we can prove that $L^{\prime}_{k}$ is in $\mathrm{DCFL}[2]$.
@@
(ii) Consider the case where $k$ is an even number with $k\geq 3$. Given a string $w$ of the form $w_{2}\#w_{4}\#\cdots\#w_{k-2}\#w_{k}\#w_{k-1}\#\cdots\#w_{3}\#w_{1}$, we succinctly write $\tilde{w}^{(-)}$ for the string $w_{4}\#\cdots\#w_{k-2}\#w_{k}\#w_{k-1}\#\cdots\#w_{3}$ so that $w$ equals $w_{2}\#\tilde{w}^{(-)}\#w_{1}$. Let $A_{(1)}=\{w\mid n_{1}=m_{1},n_{2}=m_{2},\tilde{w}^{(-)}\in L_{k-2}\}$, $A_{(2)}=\{w\mid n_{1}=m_{1},n_{2}<m_{2},\tilde{w}^{(-)}\in L^{\prime}_{k-2}\}$, $A_{(3)}=\{w\mid n_{1}<m_{1},m_{2}=p_{2},\tilde{w}^{(-)}\in L_{k-2}\}$, and $A_{(4)}=\{w\mid n_{1}<m_{1},m_{2}<p_{2},\tilde{w}^{(-)}\in L^{\prime}_{k-2}\}$, where $w$ is of the above-mentioned form. By induction hypothesis for $L_{k-2}$, $L_{k-2}\in\mathrm{DCFL}[2^{(k-2)-1}]$ implies that $A_{(j)}\in\mathrm{DCFL}[2^{(k-2)-1}]$ for every index $j\in[4]$. Since $L_{k}=A_{(1)}\cup A_{(2)}\cup A_{(3)}\cup A_{(4)}$, it instantly follows that $L_{k}\in\mathrm{DCFL}[4\cdot 2^{(k-2)-1}]=\mathrm{DCFL}[2^{k-1}]$.
(iii) The case of odd $k\geq 3$ is treated similarly to (ii).
@@
(2) As the second step, we prove that $L_{k}\notin\mathrm{DCFL}[2^{k-1}-1]$. Assuming to the contrary that $L_{k}\in\mathrm{DCFL}[2^{k-1}-1]$, we apply the first pumping lemma for $\mathrm{DCFL}[2^{k-1}-1]$ to $L_{k}$.
@@
As for the base case of $k=2$, the first pumping lemma for $\mathrm{DCFL}$ easily leads to $L_{2}\notin\mathrm{DCFL}$, from which we obtain $L_{2}\notin\mathrm{DCFL}[1]$. Next, we focus on the case of $k\geq 3$. Toward a contradiction, we assume that $L_{k}\in\mathrm{DCFL}[2^{k-1}-1]$. In the case where $k$ is even, we first define the index set $I_{k}=\{1,2,3,4\}^{k-1}\times\{1,2\}$. We then take $k$ sufficiently large positive integers $n_{1},n_{2},\ldots,n_{k}$, which are all different from each other. For each index $j\in[k]$, we set $w_{j}^{(1)}=a^{n_{j}}b^{n_{j}}c^{n_{j}-1}$, $w_{j}^{(2)}=a^{n_{j}}b^{n_{j}+1}c^{n_{j}}$, $w_{j}^{(3)}=a^{n_{j}}b^{n_{j}-1}c^{n_{j}-1}$, and $w_{j}^{(4)}=a^{n_{j}}b^{n_{j}-1}c^{n_{j}}$. For a given string $s$ in $I_{k}$ of the form $s_{2}s_{4}\cdots s_{k-2}s_{k}s_{k-1}\cdots s_{3}s_{1}$, we define $\alpha_{s}=w_{2}^{(s_{2})}\#w_{4}^{(s_{4})}\#\cdots\#w_{k-2}^{(s_{k-2})}\#w_{k}^{(s_{k})}\#w_{k-1}^{(s_{k-1})}\#\cdots\#w_{3}^{(s_{3})}\#w_{1}^{(s_{1})}$. For any $j\in[k]$, we define $\alpha_{s}^{(j)}=w_{2}^{(s_{2})}\#w_{4}^{(s_{4})}\#\cdots\#w_{j-1}^{(s_{j})}\#$ and $\beta_{s}^{(j)}=\#w_{j+1}^{(s_{j+1})}\#\cdots\#w_{1}^{(s_{1})}$ so that $\alpha_{s}=\alpha_{s}^{(j)}\beta_{s}^{(j)}$ holds.

### v3

@@ 1.2 New Separation Results
As noted in Section 1.1, some of the fundamental properties associated with $\mathrm{DCFL}(d)$ heavily rely on the single separation result (*) of Liu and Weiner. However, Liu and Weiner’s technical tool that leads to their main result does not seem to withstand a wide variety of direct applications. It is thus desirable to develop a new, simple, and practical technical tool that can find numerous applications for conducting a future study on $\mathrm{DCFL}(d)$ as well as $\mathrm{DCFL}[d]$. Our main purpose of this exposition is therefore to present a simple but powerful and practical technical tool, which is applicable to numerous languages. In a course of analyzing parse trees of (strict) dcf languages, Harrison and Havel [8] developed an iteration theorem for $\mathrm{DCFL}$. A similar approach has been taken to introduce numerous forms of pumping lemmas and iteration theorems for variants of dcf languages in, e.g., [7, 11, 12, 13, 20, 32]. Along this long line of study, this exposition also aims at developing pumping lemmas for $\mathrm{DCFL}[d]$ for any $d\geq 1$ (formulated as Lemmas 2.1 and 2.2 in Section 2.2), which may contribute to the deepening of our understandings of both $\mathrm{DCFL}[d]$ and $\mathrm{DCFL}(d)$. See Fig. 1.
@@ 2.3 Pushdown Automata in an Ideal Shape
It was shown in [26, 28] that any 1ppda can be converted into its “equivalent” 1ppda in an ideal shape. We say that two 1ppda’s are error-equivalent if, for any input $x$, their acceptance probabilities coincide.
@@ 3 New Technical Tools: Pumping Lemmas for DCFL[$d$]
In the past literature, numerous pumping lemmas (as well as iteration theorems) have been proposed. Interestingly, the form of these pumping lemmas varies. For instance, the iteration theorem of Harrison and Havel [8, Theorem 2.3] for $\mathrm{DCFL}$ focuses on a single input string in a target language, whereas Yu’s pumping lemma for $\mathrm{DCFL}$ [32, Lemma 1] deals with two input strings in the language. We wish to take the latter approach in this exposition.
@@
As a special case of $d=1$ in Lemma 3.1, we instantly obtain Yu’s pumping lemma for $\mathrm{DCFL}$ [32, Lemma 1], which was proven by a grammar-based argument. By sharp contrast, a few machine-based arguments were employed in the past literature in order to prove generic properties of $\mathrm{CFL}$ and $\mathrm{DCFL}$. As a recent example, the swapping lemma for context-free languages was proven in [23, Lemma 4.1] and [25, Corollary 4.2] by exploring various properties of one-way nondeterministic pushdown automata (or 1npda’s). Such machine-based arguments as well as various properties of machines that these arguments have brought in are not only interesting in its own right but also quite useful to analyze families of languages recognized by different types of machines, such as one-way depth-bounded storage automata [29, 30, 31], which naturally expand 1npda’s and 1dpda’s but no grammar characterizations of them are currently known to exploit.
@@ 4 Proofs of Three Separation Claims
(1) It is well-known that $Pal$ is in $\mathrm{CFL}$. Hence, we wish to verify that $Pal$ is also in $\mathrm{co}\mbox{-}\mathrm{CFL}$. Let $ODD=\{w\in\{0,1\}^{*}\mid|w|\text{ is odd }\}$. Note that $\overline{Pal}$ coincides with $\{xy\mid|x|=|y|,x,y\in\{0,1\}^{*},y\neq x^{R}\}\cup ODD$. Consider the following 1npda. Given an input $w$, choose either 0 or $1$ nondeterministically. If 0 is chosen, then we check whether $w\in ODD$. Otherwise, we split $w$ nondeterministically into $xy$, store $x$ into a stack, pop $x$ symbol by symbol while reading $y$, and check that $y=x^{R}$. If $|x|\neq|y|$, then we reject $w$. Assume otherwise. If we discover any discrepancy between $y$ and $x^{R}$, then we accept $w$; otherwise, we reject it. It is obvious that this machine correctly recognizes $\overline{Pal}$, and thus $Pal$ belongs to $\mathrm{co}\mbox{-}\mathrm{CFL}$.
@@
In the case where the condition (1) fails, let us consider three factorizations $x^{\prime}=x_{1}x_{2}x_{3}$, $y=y_{1}y_{2}y_{3}$, and $z=z_{1}z_{2}z_{3}$ with $|x_{2}|\geq 1$ and $|x_{2}x_{3}|\leq c$ satisfying $x_{1}x_{2}^{i}x_{3}y_{1}y_{2}^{i}y_{3},x_{1}x_{2}^{i}x_{3}z_{1}z_{2}^{i}z_{3}\in Pal$ for any number $i\geq 0$. We conclude that $x_{2}$ is in the $(2j_{2}+2d+1)$th block of $x^{\prime}$ because of $|x_{2}x_{3}|\leq c$. By taking $i=0$, the size of this block becomes less than $n$. Since this block is located in the right side of the center of $x^{\prime}y$, we obtain $x_{1}x_{3}y_{1}y_{3}\notin Pal$, a contradiction.
@@
The second pumping lemma then ensures the existence of two indices $j_{1},j_{2}\in[d+1]$ with $j_{1}<j_{2}$ such that $(h_{j_{1}},h_{j_{2}})$ satisfies one of the conditions (1)–(5) of the lemma. We then define $x^{\prime}=w_{1}\#\cdots w_{d+1}\#^{2}v_{1}\#\cdots\#(v_{j_{1}})^{(-)}_{\text{first}}$, $y=(\hat{v}_{j_{1}})^{(-)}_{\text{second}}\#v_{j_{1}+1}\#\cdots\#v_{d+1}$, and $z=(v_{j_{1}})^{(-)}_{\text{second}}\#v_{j_{1}+1}\#\cdots\#v_{j_{2}-1}\#\hat{v}_{j_{2}}\#v_{j_{2}+1}\#\cdots\#v_{d+1}$. Note that $h_{j_{1}}=x^{\prime}y$ and $h_{j_{2}}=x^{\prime}z$. By the above statement (*), the conditions (1)–(4) do not hold. We thus look into the condition (5). A nondegenerate iterative pair $(x_{4}x_{5},x_{7}u_{1})$ of $x^{\prime}y$ for $L^{\prime\prime}$ in the condition (5) must satisfy that $x_{4}x_{5}\sqsubseteq w_{j_{1}}$ and $x_{7}u_{1}\sqsubseteq\hat{v}_{j_{1}}$. Similarly, if a nondegenerate iterative pair $(x_{3}x_{4},u^{\prime}_{2})$ of $x^{\prime}z$ for $L^{\prime\prime}$ exists, then we obtain $x_{3}x_{4}\sqsubseteq w_{j_{2}}$ and $u^{\prime}_{2}\sqsubseteq\hat{v}_{j_{2}}$. However, these contradict the assumption of $j_{1}<j_{2}$. The case of $(x_{2},u^{\prime}_{2})$ in the condition (5) is similarly treated. Therefore, $L^{\prime\prime}$ is not in $\mathrm{DCFL}[d]$, as requested. This completes the proof. ∎
@@
Proposition 1.5. (rephrased) For any $k\geq 2$, $k\mbox{-}\mathrm{LDA}\cap\mathrm{DCFL}[2^{\lfloor k/2\rfloor}]\nsubseteq(k-1)\mbox{-}\mathrm{LDA}\cup\mathrm{DCFL}[2^{\lfloor k/2\rfloor}-1]$.
@@
In this proof, we intend to use a series of special languages, each of which is a slight modification of the language used in [26, arXiv version], introduced directly from Hibbard’s [9, Section 4] example language, and is shown to be in $k\mbox{-}\mathrm{LDA}$ but not in $(k-1)\mbox{-}\mathrm{LDA}$ for each $k\geq 2$ [9, 26]. We call this new language by $L_{k}$.
Let $k\geq 2$ be any integer. For each fixed index $i\in[k]$, we write $w_{i}$ for an arbitrary string of the form $a^{n_{i}}b^{m_{i}}c^{p_{i}}$ for any numbers $n_{i},m_{i},p_{i}\in\mathbb{N}$. The language $L_{k}$ is composed of all strings $w$ of the form $w_{2}\#w_{4}\#\cdots\#w_{k-2}\#w_{k}\#w_{k-1}\#\cdots\#w_{5}\#w_{3}\#w_{1}$ if $k$ is even, and $w_{2}\#w_{4}\#\cdots\#w_{k-1}\#w_{k}\#w_{k-2}\#\cdots\#w_{5}\#w_{3}\#w_{1}$ if $k$ is odd, together with the following conditions (a)–(e).
@@
    It holds that $n_{1}\leq m_{1}\leq n_{1}+1$.
@@
    If $n_{2}\leq m_{2}\leq n_{2}+1$, then $n_{1}=m_{1}$ holds. If $m_{2}\leq p_{2}\leq m_{2}+1$, then $m_{1}=n_{1}+1$ holds.
  4. (d)
    Let $j$ denote any index in $3,k]_{\mathbb{Z}}$. If $n_{j}\leq m_{j}\leq n_{j}+1$, then either $n_{j-1}=m_{j-1}$ or $m_{j-1}=p_{j-1}$ holds. If $m_{j}\leq p_{j}\leq m_{j}+1$, then either $m_{j-1}=n_{j-1}+1$ or $p_{j-1}=m_{j-1}+1$ holds.
  5. (e)
@@
For instance, $L_{2}$ equals $\{a^{n_{2}}b^{m_{2}}c^{p_{2}}\#a^{n_{1}}b^{m_{1}}c^{p_{1}}\mid(n_{1}=m_{1}\wedge n_{2}=m_{2})\text{ or }(m_{1}=n_{1}+1\wedge m_{2}=p_{2})\}$. For convenience, we also define another language $L^{\prime}_{k}$ obtained from $L_{k}$ by replacing the last term $w_{1}=a^{n_{1}}b^{m_{1}}c^{p_{1}}$ of $L_{k}$ with $a^{p_{1}}b^{m_{1}}c^{n_{1}}$ but we treat $(p_{1},m_{1},n_{1})$ as $(n_{1},m_{1},p_{1})$ in (a)–(e).
@@
Following [26, arXiv version], $L_{k}$ can be proven to belong to $k\mbox{-}\mathrm{LDA}$ but it is not in $(k-1)\mbox{-}\mathrm{LDA}$. It thus suffices to show by induction on $k$ that $L_{k}\in\mathrm{DCFL}[2^{\lfloor k/2\rfloor}]$ and $L_{k}\notin\mathrm{DCFL}[2^{\lfloor k/2\rfloor}-1]$. To do so, we want to take the following two steps (1)–(2). For readability, we assume in (1)–(2) that $n_{i}$, $m_{i}$, and $p_{i}$ express numbers in $\mathbb{N}$ for all $i$’s.
@@
(1) As the first step, we prove the membership of $L_{k}$ to $\mathrm{DCFL}[2^{\lfloor k/2\rfloor}]$.
@@
(i) When $k=2$, we define $A_{(1)}=\{a^{n_{2}}b^{m_{2}}c^{p_{2}}\#a^{n_{1}}b^{m_{1}}c^{p_{1}}\mid n_{1}=m_{1},n_{2}=m_{2}\}$ and $A_{(2)}=\{a^{n_{2}}b^{m_{2}}c^{p_{2}}\#a^{n_{1}}b^{m_{1}}c^{p_{1}}\mid m_{1}=n_{1}+1,m_{2}=p_{2}\}$. Clearly, $L_{2}$ equals $A_{(1)}\cup A_{(2)}$, and thus $L_{2}$ belongs to $\mathrm{DCFL}[2]$. Similarly, we can prove that $L^{\prime}_{2}$ is in $\mathrm{DCFL}[2]$.
(ii) In the case of $k=3$, we set $A_{(1)}=\{w\mid n_{1}=m_{1},(n_{2}=m_{2}\wedge n_{3}=m_{3})\text{ or }(m_{2}=n_{2}+1\wedge m_{3}=p_{3})\}$ and $A_{(2)}=\{w\mid m_{1}=n_{1}+1,(m_{2}=p_{2}\wedge n_{3}=m_{3})\text{ or }(p_{2}=m_{2}+1\wedge m_{3}=p_{3})\}$. It follows that $L_{3}=A_{(1)}\cup A_{(2)}\in\mathrm{DCFL}[2]$. A similar argument proves that $L^{\prime}_{3}$ is in $\mathrm{DCFL}[2]$.
@@
(iii) Consider the case where $k$ is an even number with $k\geq 4$. Given a string $w$ of the form $w_{2}\#w_{4}\#\cdots\#w_{k-2}\#w_{k}\#w_{k-1}\#\cdots\#w_{3}\#w_{1}$, we succinctly write $\tilde{w}^{(-)}$ for the string $w_{4}\#\cdots\#w_{k-2}\#w_{k}\#w_{k-1}\#\cdots\#w_{3}$ so that $w$ equals $w_{2}\#\tilde{w}^{(-)}\#w_{1}$. Let $A_{(1)}=\{w\mid(n_{1}=m_{1}\wedge n_{2}=m_{2}\wedge\tilde{w}^{(-)}\in L_{k-2})\text{ or }(n_{1}=m_{1}\wedge m_{2}=n_{2}+1\wedge\tilde{w}^{(-)}\in L^{\prime}_{k-2})\}$ and $A_{(2)}=\{w\mid(m_{1}=n_{1}+1\wedge m_{2}=p_{2}\wedge\tilde{w}^{(-)}\in L_{k-2})\text{ or }(m_{1}=n_{1}+1\wedge p_{2}=m_{2}+1\wedge\tilde{w}^{(-)}\in L^{\prime}_{k-2})\}$, where $w$ is of the above-mentioned form. By induction hypothesis, $L_{k-2},L^{\prime}_{k-2}\in\mathrm{DCFL}[2^{\lfloor(k-2)/2\rfloor}]$ implies that $A_{(j)}\in\mathrm{DCFL}[2^{\lfloor(k-2)/2\rfloor}]$ for every index $j\in[2]$. Since $L_{k}=A_{(1)}\cup A_{(2)}$, it instantly follows that $L_{k}\in\mathrm{DCFL}[2\cdot 2^{\lfloor(k-2)/2\rfloor}]=\mathrm{DCFL}[2^{\lfloor k/2\rfloor}]$. A similar argument proves that $L^{\prime}_{k}\in\mathrm{DCFL}[2^{\lfloor k/2\rfloor}]$.
(iv) The case of odd $k\geq 3$ is treated similarly to (iii).
@@
(2) As the second step, we prove that $L_{k}\notin\mathrm{DCFL}[2^{\lfloor k/2\rfloor}-1]$. Assuming to the contrary that $L_{k}\in\mathrm{DCFL}[2^{\lfloor k/2\rfloor}-1]$, we apply the pumping lemmas for $\mathrm{DCFL}[2^{\lfloor k/2\rfloor}-1]$ to $L_{k}$.
@@
As for the base case of $k=2$, the first pumping lemma for $\mathrm{DCFL}$ easily leads to $L_{2}\notin\mathrm{DCFL}$, from which we obtain $L_{2}\notin\mathrm{DCFL}[1]$. Next, we focus on the case of $k\geq 3$. Toward a contradiction, we assume that $L_{k}\in\mathrm{DCFL}[2^{\lfloor k/2\rfloor}-1]$. In the case where $k$ is even, we first define the index set $I_{k}=\{1,2,3,4\}^{k-1}\times\{1,2\}$. We then take $k$ sufficiently large (and thus larger than $c$) positive integers $n_{1},n_{2},\ldots,n_{k}$, which are all different from each other. For each index $j\in[k]$, we set $w_{j}^{(1)}=a^{n_{j}}b^{n_{j}}c^{n_{j}-1}$, $w_{j}^{(2)}=a^{n_{j}}b^{n_{j}+1}c^{n_{j}}$, $w_{j}^{(3)}=a^{n_{j}}b^{n_{j}-1}c^{n_{j}-1}$, and $w_{j}^{(4)}=a^{n_{j}}b^{n_{j}-1}c^{n_{j}}$. For a given string $s$ in $I_{k}$ of the form $s_{2}s_{4}\cdots s_{k-2}s_{k}s_{k-1}\cdots s_{3}s_{1}$, we define $\alpha_{s}=w_{2}^{(s_{2})}\#w_{4}^{(s_{4})}\#\cdots\#w_{k-2}^{(s_{k-2})}\#w_{k}^{(s_{k})}\#w_{k-1}^{(s_{k-1})}\#\cdots\#w_{3}^{(s_{3})}\#w_{1}^{(s_{1})}$. In what follows, we concentrate on such a series $s$ that makes $\alpha_{s}$ belong to $L_{k}$. For the language $L_{k}$, let us recall the $2^{\lfloor k/2\rfloor}$ dcf languages, say, $A_{1},A_{2},\ldots,A_{2^{\lfloor k/2\rfloor}}$, which have been constructed recursively by (i)–(iv). For each $A_{i}$, we choose the smallest $s$ from $I_{k}$ such that $\alpha_{s}$ falls in $A_{i}$, where $s$ is viewed as an integer vector of dimension $k$. We then form a set of those minimum vectors $s$ and call it by $A_{vec}$. For instance, when $k=3$, $A_{vec}$ equals $\{111,312\}$; when $k=4$, $A_{vec}$ is $\{1111,1321,3112,3322\}$; when $k=5$, $A_{vec}$ matches $\{11111,13121,31112,33122\}$.